Use the method of disks/rings to determine the volume of the solid obtained by rotating the region bounded by y=7-x^2, x=-2, x=2 and the x-axis about the x-axis.


1
Expert's answer
2021-04-15T07:52:32-0400

A graph of the function "y=7-x^2" between "x=-2" and "x=2" has the form:


In case we rotate the plot around "x"-axis and then fix "x", we will get a circle of the radius "R=7-x^2". The area enclosed by it is "S=\pi R^2=\pi(7-x^2)^2". In order to find the volume of the solid, we consider equal intervals "\Delta x_i=x_{i}-x_{i-1}, i=1,...,n", where "-2=x_0<x_1<x_2<...<x_n=2" and the following sum: "S_n=\sum_{i=0}^n\pi(7-x_i^2)^2\Delta x_i". In case we take the limit "\Delta x_i\rightarrow0" (it is equivalent to "n\rightarrow\infty" ), we will receive the integral that provides the volume of the solid: "V=\pi\int_{-2}^2(7-x^2)^2dx=\pi(\int_{-2}^2(49-14x^2+x^4)dx)=\pi(49x-\frac{14}{3}x^3+\frac{1}{5}x^5)|_{-2}^2=\pi(196-\frac{14\cdot16}{3}+\frac{64}{5})=\frac{2012}{15}\pi"


Answer: The volume of the solid is "\frac{2012}{15}\pi"